OpenStudy (anonymous):

Part 1: Provide the solution to the equation 25 = square root of 5 (2x + 6). (3 points) Part 2: Using complete sentences, explain the strategy used in solving this equation the way you did. (3 points) please help im so lost...

OpenStudy (mathstudent55):

This? \( 25 = \sqrt{5(2x + 8)} \)

OpenStudy (anonymous):

only the 5 with the exponent of 2x+8 @mathstudent55

OpenStudy (mathstudent55):

This? \(25 = \sqrt{5^{2x + 8}} \)

OpenStudy (anonymous):

25 \[\sqrt{5}^{2x+8}\]

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

\( 25 = \sqrt{5}^{2x + 8} \) Rewrite 5 as \(5^2\). Also, \( \sqrt{5} = 5^{\frac{1}{2}} \).

OpenStudy (mathstudent55):

\( 5^{2} = 5^{ \frac{1}{2}(2x +8)} \) \( 5^{2} = 5^{ x + 4} \) That means: \(2 = x + 4\)

OpenStudy (anonymous):

thanks sooo much @mathstudent55